### v1

# Volumetric Functional Maps

Abstract.
The computation of volumetric correspondences between 3D shapes has great potential for medical and industrial applications. In this work, we pave the way for spectral volume mapping, extending for the first time the functional maps framework from the surface setting to the volumetric domain. We show that the eigenfunctions of the volumetric Laplace operator define a functional space that is suitable for high-quality signal transfer. We also experiment with various techniques that edit this functional space, porting them from the surface to the volume setting. We validate our method on novel volumetric datasets and on tetrahedralizations of well established surface datasets, also showcasing practical applications involving both discrete and continuous signal mapping, for segmentation transfer, mesh connectivity transfer and solid texturing. Last but not least, we show that considering the volumetric spectrum greatly improves the accuracy for classical shape matching tasks among surfaces, consistently outperforming existing surface-only spectral methods.
Functional maps; volume mapping; shape matching; shape correspondence
submissionid: 598ccs: Computing methodologies Shape analysisccs: Computing methodologies Volumetric modelsccs: Computing methodologies Computer graphics
Figure 1. A visual representation of our pipeline. The eigenfunctions of the Laplace-Beltrami operator for tetrahedral meshes (left) are used to compute a functional map among volumes (middle). The basis alignment can be applied for several tasks, such as computation of volumetric correspondences, generation of piece-wise linear simplicial maps, and volumetric segmentation transfer (right).
1. Introduction
The computation of volumetric correspondences between solid objects is becoming more and more important for applications. To make some concrete examples, volume mappings enable non-invasive medical diagnosis, by warping the digital copy of a human organ into its canonical position without extracting it from the patient’s body (Abulnaga et al., 2021). They are also important for hexmeshing (Pietroni et al., 2022), statistical shape analysis (Huang et al., 2019) and for Industry 4.0, where physical simulations computed on a predictive digital twin must be aligned with the other digital views of the same real environment for optimized decision making and failure avoidance (Rasheed et al., 2020).
Maps between 3D objects are often represented in a computer either by exploiting one-to-one per vertex correspondences between two simplicial meshes with same connectivity, defining a so called piece-wise map (Section 2.1), or putting into correspondence two sets of basis functions that allow to express the same signal in two alternative domains, as done by spectral approaches for surface meshes (Ovsjanikov et al., 2012). Both piece-wise and spectral approaches have been extensively studied in the surface setting and can nowadays be considered rather mature, as proved by the numerous academic and commercial tools that incorporate these technologies, such as Blender, Maya and 3DS Max. Conversely, the equivalent task of computing volumetric correspondences between two solid objects largely remains a scientific open problem (Pietroni et al., 2022, §8.2).
In recent years various attempts have been made to extend piece-wise linear approaches to the volume setting. This has proved to be an extremely challenging task, for which no fully robust and efficient methods have been devised yet (Section 2.1). Conversely, to our knowledge no attempt has been made so far to extend functional mapping from the surface to the volume setting (Section 2.2).
In this article, we propose this extension for the first time and show that, conversely from the piece-wise linear setting, functional mapping naturally scales to volumes with minimal effort, providing a theoretically sound and practically useful platform for the computation of volumetric correspondences between solid shapes.
Using the spectrum of the volumetric Laplacian we define a functional space that is completely agnostic w.r.t. the underlying discrete mesh and we use it to create correspondences between solid meshes with different density and connectivity. Interestingly, in the paper we show that many recent developments used to improve efficiency by reducing the size of the functional space without sacrificing accuracy (Melzi et al., 2019b; Maggioli et al., 2021) can also be exploited in the volume setting, permitting to find an optimal balance between map quality and performances.
We validate our volumetric pipeline (showcased in Figure 1) on a variety of datasets, including tetrahedralizations of prominent reference benchmarks widely used in the functional mapping literature. Besides measuring the accuracy and performances of our tool on ground truth data (Section 5), we showcase a variety of practical applications, including connectivity transfer of a given tetrahedral mesh to a target domain to generate a piece-wise linear map (Figure 6), segmentation transfer from a statistical medical shape to acquired organs (Figure 9) and solid texturing (Figure 5).
All in all, this article opens for the first time to the computation of volumetric correspondences with spectral approaches. To this end, we believe our contributions will foster new research in this direction, positioning volume spectral mapping as a flexible and practical alternative to existing piece-wise linear approaches.
2. Related Works
Here, we discuss the main approaches for the computation of volume maps with the piece-wise linear paradigm and surface correspondence with the functional maps framework. In doing this, we highlight the existing attempts to scale from the surface to the volume setting (or lack thereof) for each class of algorithms.
2.1. Piece-wise Maps
Piece-wise maps define correspondences between two simplicial meshes sharing the same amount of vertices and same connectivity. Starting from explicit one-to-one per vertex correspondences the map is implicitly extended inside mesh elements by means of linear barycentric interpolation. This operation is uniquely defined for simplices in any dimension (Hormann, 2014) and can therefore be used for both surface (triangle) and volume (tetrahedral) meshes.
A variety of algorithms ensure the ability to generate piece-wise linear surface maps to elementary domains such as triangles (Finnendahl et al., 2023), convex polygons (Tutte, 1963; Floater, 1997; Livesu, 2024b; Shen et al., 2019), star-shaped polygons (Livesu, 2024a) and spheres (Gotsman et al., 2003; Praun and Hoppe, 2003). Mappings to these basic shapes (or to assemblies of them) are the main ingredient to generate higher level maps between general shapes, which are typically obtained via composition, overlaying two surface maps one onto the other to define inter-surface correspondences (Schmidt et al., 2019; Schreiner et al., 2004; Weber and Zorin, 2014; Lee et al., 1998; Khodakovsky et al., 2003; Praun et al., 2001; Schmidt et al., 2020; Livesu, 2021; Garner et al., 2005; Shi et al., 2016; Aigerman and Lipman, 2016; Li et al., 2008; Schmidt et al., 2023; Kraevoy and Sheffer, 2004).
Lifting the same idea to the volume setting is possible in principle but extremely challenging in practice (Livesu, 2020a; Cherchi and Livesu, 2023; Nigolian et al., 2024). Even constructing an injective map to a basic convex polyhedron has proved to be a surprisingly difficult task. Surface algorithms such as the Tutte embedding are known to fail on general volume meshes (Alexa, 2023; Chilakamarri et al., 1995). More recent methods, such as (Livesu, 2024b, a; Finnendahl et al., 2023) may potentially scale to volumes but concrete extensions do not exist yet. To date, the only existing algorithms for robust volume embedding are (Nigolian et al., 2023, 2024; Hinderink and Campen, 2023; Campen et al., 2016). All of them are extremely inefficient both in terms of running times and memory consumption, mainly due to the need of using exact numerical constructions and exponential mesh refinement. They are therefore unsuitable for practical applications.
A less robust but more practical alternative to the map composition pipeline consists in initializing a non injective volume map directly, with Tutte 3D or some other non robust method, and then restore the correct orientation for elements that flip or vanish through the map via vertex relocation. In the volume setting orientation (injectivity) constraints translate to sign constraints on the determinant of the $3\times 3$ Jacobian matrix associated to each per element transformation, leading to an optimization problem subject to cubic constraints which is numerically intractable. Most of the research in the field focuses on how to formulate (and relax) the orientation constraints to make the problem tractable. Removing inverted elements from a given mesh is also a well known problem in computer simulation, where it is often referred to as mesh untangling (Knupp, 2001; Pietroni et al., 2022). Note that in the general case mesh untangling is an ill-posed problem, because it may not admit a valid solution for a fixed mesh connectivity. Additional degrees of freedom must be inserted via mesh refinement to open the feasible region (see (Livesu, 2020b, Fig.1) for a 2D example). Existing volume approaches employ minimization of flip-preventing energies (Du et al., 2020; Garanzha et al., 2021; Abulnaga et al., 2023), operate on convex subsets of the feasible region (Kovalsky et al., 2014) or project onto the space of bounded distortion mappings (Aigerman and Lipman, 2013). Due to the non linearity of the problem and the non existence of a feasible starting solution, none of these methods guarantees the correctness of the result. Multiple failure cases have been reported by testing the most prominent existing algorithms on available datasets (Nigolian et al., 2023, §6.3).
2.2. Functional Maps
One of the most fundamental topics in shape analysis is the shape matching task. Given a pair of discretized surfaces of 3D objects, shape matching estimates a meaningful pointwise correspondence between them. If the shapes are rigid, we refer to this task as rigid shape matching. Otherwise, we refer to it as deformable shape matching or, for brevity, shape matching, which is more general and usually more challenging to solve. Many applications are related to this task, from statistical shape analysis (Bogo et al., 2014) and medical imaging (Magnet et al., 2023) to deformation transfer (Sumner and Popović, 2004) among many others. For this reason, many researchers target this task (Sahillioğlu, 2020; Deng et al., 2022). Among the different approaches, the functional maps framework (Ovsjanikov et al., 2012, 2016) has attracted significant attention. This framework proposes to solve the shape matching problem by estimating a correspondence between functions defined on the surfaces instead of focusing on correspondence among points which are hard to optimize. Furthermore, exploiting a basis for the space of functions defined on the surfaces, for which a standard choice arises from the extension of Fourier theory to non-Euclidean domains (Taubin, 1995; Lévy and Zhang, 2010), the correspondence among functions can be compactly encoded in a matrix with dimensions equal to the size of the selected basis. Following this direction, a variety of works tried to improve the framework by defining additional functional constraints (Nogneng and Ovsjanikov, 2017; Ren et al., 2018; Donati et al., 2022), proposing alternative bases (Melzi et al., 2018; Nogneng et al., 2018; Maggioli et al., 2021; Marin et al., 2019, 2020), or defining an alternative procedure to extract the correspondence from a given functional map (Rodola et al., 2017; Melzi et al., 2019b; Ren et al., 2020, 2021; Ezuz and Ben-Chen, 2017; Pai et al., 2021). Recently, by proposing two alternatives for scaling the functional maps approach to high-resolution meshes (Magnet and Ovsjanikov, 2023; Maggioli et al., 2025) addressed the issue of scalability of the functional maps framework. The functional maps framework additionally gives rise to a family of data-driven approaches that exploit machine learning techniques to solve the shape matching problem exploiting the functional representation as a prior (Litany et al., 2017; Donati et al., 2020; Sharp et al., 2022; Cao et al., 2024, 2023; Attaiki and Ovsjanikov, 2024). The ones listed above are just a subset of the literature related to the functional maps framework. Its impact in computer graphics and related fields has been impressive, and many applications have been addressed exploring this technique, from statistical shape analysis in medicine (Maccarone et al., 2024) to shape deformation (Sundararaman et al., 2024). Despite its success, to the best of our knowledge, this framework has never been applied to volumetric data, leaving its potential impact entirely unexplored in volumetric data processing and related tasks.
3. Background and Notation
We discretize volumes as tetrahedral meshes $\mathcal{M}=(V_{\mathcal{M}},T_{\mathcal{M}})$, where:
  • $V_{\mathcal{M}}\subset\mathbb{R}^{3}$ is a set of vertices in 3D space;
  • $T_{\mathcal{M}}\subset V_{\mathcal{M}}^{4}$ is a set of tetrahedra connecting vertices in $V_{\mathcal{M}}$.
The inset figure depicts a section of a tetrahedral mesh, with the original external surface shown in transparency.
Scalar functions $f:V_{\mathcal{M}}\to\mathbb{R}$ over a volume mesh $\mathcal{M}$ are discretized as vectors $\bm{f}\in\mathbb{R}^{|V_{\mathcal{M}}|}$. Given a scalar function $f:V_{\mathcal{M}}\to\mathbb{R}$ over a volume mesh $\mathcal{M}$, its restriction $f_{\partial}:V_{\partial\mathcal{M}}\to\mathbb{R}$ to the external surface $\partial\mathcal{M}$ is discretized as a vector $\bm{f}_{\partial}\in\mathbb{R}^{|V_{\partial\mathcal{M}}|}$.
We refer to the external surface of $\mathcal{M}$ as the triangular mesh $\partial\mathcal{M}=(V_{\partial\mathcal{M}},F_{\partial\mathcal{M}})$, where $V_{\partial\mathcal{M}}\subset V_{\mathcal{M}}$is the set of surface vertices of $\mathcal{M}$; $F_{\partial\mathcal{M}}\subset V_{\partial\mathcal{M}}^{3}$is the set of surface triangles defined by the tetrahedra in $T_{\mathcal{M}}$.
Given a tetrahedral (or surface) mesh $\mathcal{M}$, and given two vertices $v_{i},v_{j}\in\mathcal{M}$, we refer to the geodesic distance between $v_{i}$ and $v_{j}$ over $\mathcal{M}$ as $d_{\mathcal{M}}(v_{i},v_{j})$. There exist several algorithms to estimate such a distance on discrete meshes (Crane et al., 2020). For the sake of simplicity, in the remainder of the paper we will always measure $d_{\mathcal{M}}(v_{i},v_{j})$ using Dijkstra’s algoritm (Dijkstra, 1959) on the edge connectivity of $\mathcal{M}$.
Shape Matching and Correspondences
Given two surface meshes $\mathcal{M}=(V_{\mathcal{M}},F_{\mathcal{M}})$ and $\mathcal{N}=(V_{\mathcal{N}},F_{\mathcal{N}})$, the problem of shape matching is the problem of finding a correspondence $\pi:V_{\mathcal{M}}\to V_{\mathcal{N}}$ that maps vertices of $\mathcal{M}$ into vertices of $\mathcal{N}$ (Loncaric, 1998). In most cases, the correspondence $\pi$ has to satisfy semantic constraints or has to be an isometry (Sahillioğlu, 2020; Deng et al., 2022). An isometry is a correspondence $\pi$ that preserves the geodesic distances between pairs of vertices, namely
(1) $\forall v_{i},v_{j}\in V_{\mathcal{M}},\quad d_{\mathcal{M}}(v_{i},v_{j})=d_{\mathcal{N}}(\pi(v_{i}),\pi(v_{j}))\,.$
Functional Maps
The problem of finding an isometric correspondence, in its general setting, is NP-hard (Benkner et al., 2021), and thus many solutions have been developed for computing an approximation. Among these, the functional maps approach has grown popular due to its high efficiency and the smoothness of the resulting correspondences (Ovsjanikov et al., 2012).
Instead of searching directly for a correspondence $\pi$ between two surfaces $\mathcal{M}$ and $\mathcal{N}$, the functional maps approach aims at finding a linear map $T_{\pi}:\mathcal{F}(\mathcal{N})\to\mathcal{F}(\mathcal{M})$, induced by $\pi$, from the space $\mathcal{F}(\mathcal{N})$ of real-valued functions over $\mathcal{N}$ to the space $\mathcal{F}(\mathcal{M})$ of real-valued functions over $\mathcal{M}$. Given any function $f:V_{\mathcal{N}}\to\mathbb{R}$, the map $T_{\pi}$ gives the corresponding function $T_{\pi}(f):V_{\mathcal{M}}\to\mathbb{R}$ that maps $v_{i}\mapsto f(\pi(x))$, for each $v_{i}\in V_{\mathcal{M}}$.
Given two sets of (orthogonal) basis functions $\mathbf{\Phi}_{\mathcal{M}},\mathbf{\Phi}_{\mathcal{N}}$, respectively for the functional space $\mathcal{F}(\mathcal{M}),\mathcal{F}(\mathcal{N})$, and a function $f:V_{\mathcal{N}}\to\mathbb{R}$, the function $f$ can be represented as a linear combination of the basis functions $\mathbf{\Phi}_{\mathcal{N}}$ as
(2) $f=\sum_{i=0}^{\infty}\alpha_{i}\phi_{\mathcal{N}}^{(i)}=\sum_{i=0}^{\infty}\langle f,\phi_{\mathcal{N}}^{(i)}\rangle_{\mathcal{N}}\phi_{\mathcal{N}}^{(i)}\,,$
where $\phi_{\mathcal{N}}^{(i)}$ is the $i$-th basis function in $\mathbf{\Phi}_{\mathcal{N}}$ and $\langle\cdot,\cdot\rangle_{\mathcal{N}}$ denotes the inner product between functions over $\mathcal{N}$. The coefficients $\alpha_{i}$ thus denotes the projection of $f$ onto the basis functions $\phi_{\mathcal{N}}^{(i)}$. A usual choice for the bases $\mathbf{\Phi}_{\mathcal{M}},\mathbf{\Phi}_{\mathcal{N}}$ are the eigenfunctions of the Laplace-Beltrami operator on $\mathcal{M}$ and $\mathcal{N}$, respectively.
Since $T_{\pi}$ is linear, the mapping of $f$ onto $\mathcal{M}$ can then be computed as
(3) $T_{\pi}(f)=\sum_{i=0}^{\infty}\alpha_{i}T_{\pi}(\phi_{\mathcal{N}}^{(i)})=\sum_{i,j=0}^{\infty}\alpha_{i}c_{i,j}\phi_{\mathcal{M}}^{(j)}\,,$
where $c_{i,j}=\langle T_{\pi}(\phi_{\mathcal{N}}^{(i)}),\phi_{\mathcal{M}}^{(j)}\rangle_{\mathcal{M}}$. By truncating the two bases to $k$ functions, the coefficients $c_{i,j}$ can be stored into a matrix $\mathbf{C}\in\mathbb{R}^{k\times k}$, leading to the matrix equation
(4) $T_{\pi}(f)\approx\mathbf{\Phi}_{\mathcal{M}}\mathbf{C}\mathbf{\Phi}_{\mathcal{N}}^{\dagger}\bm{f}\,,$
where denotes the Moore-Penrose pseudo-inverse.
Notably, $T_{\pi}(\mathbf{\Phi}_{\mathcal{N}})\approx\mathbf{\Phi}_{\mathcal{M}}\mathbf{C}$, and thus by knowing the matrix $\mathbf{C}$ it is possible to extract the unknown correspondence $\pi$ via a nearest-neighbor search $\mathrm{NN}(\mathbf{\Phi}_{\mathcal{N}},\mathbf{\Phi}_{\mathcal{M}}\mathbf{C})$.
4. Method
The functional maps framework is not bound to the domain of 2-dimensional surfaces. Indeed, the Laplace-Beltrami Operator (LBO) can be defined on Riemannian manifolds of arbitrary dimension, and it always admits a spectral decomposition. In this work, we discretize the LBO of a tetrahedral mesh $\mathcal{M}=(V_{\mathcal{M}},T_{\mathcal{M}})$ using the $n$-dimensional cotangent formula (Liao et al., 2009; Crane, 2019), which defines the discrete LBO as a pair of matrices $\mathbf{S}$ and $\mathbf{W}$, respectively denoted as stiffness and mass matrices. The spectral decomposition of the LBO can therefore be obtained by solving the generalized eigenproblem $\mathbf{S}\mathbf{\Phi}=\mathbf{W}\mathbf{\Phi}\mathbf{\Lambda}$. However, most of the properties related to the LBO eigenbasis that are regularly exploited by functional maps algorithms cannot be taken for granted due to the structural differences between tetrahedral and triangular meshes.
4.1. Functional Maps for Tetrahedral Meshes
Most successful algorithms for the estimation of the functional map between shapes rely on the property of intrinsic structures and processes to behave similarly on similar surfaces. For instance, the invariance of the LBO basis and the heat diffusion under isometric and quasi-isometric deformations plays a key role in stable functional map estimation methods (Nogneng and Ovsjanikov, 2017; Nogneng et al., 2018; Ovsjanikov et al., 2016).
LBO Eigenbasis.
When dealing with continuous Riemannian manifolds, the LBO and its eigenbasis are invariant to isometric deformations. However, in the discrete setting the position and connectivity of the vertices have a significant impact on the equivalence of the operator across different meshes, and changing the triangulation of a surface or its vertex density can strongly affect the spectral decomposition (Lescoat et al., 2020).
Moving from discrete surfaces to discrete volumes makes this problem even more evident. In a regular triangular mesh, each vertex has on average 6 neighboring vertices and 6 neighboring triangles, while in regular tetrahedral meshes the average size of a vertex neighborhood is 12 vertices and 20 tetrahedra.
Figure 2. Top: the first 5 non-constant LBO eigenfunctions on two humanoid non-isometric shapes. Bottom: eigenvalue errors distribution for the volumetric LBO (blue) and the surface LBO (orange).
To ensure that the LBO has a consistent behavior across isometric or similar volumetric shapes, we compute the spectral decomposition of the LBO on 40 pairs of shapes with identical connectivity released by Su et al. (2019) and compare both the eigenvalues and the eigenfunctions of the LBO. Our results are summarized in Figure 2. The first two rows of the figure show that two humanoid shapes share the same LBO eigenfunctions (up to sign flips), even if they are not isometric and not representing the same subject: the overall similarity is sufficient to generate the same low-frequency spectrum. To compare the eigenvalues, we first compute the vectors $\bm{\lambda}_{\mathcal{M}},\bm{\lambda}_{\mathcal{N}}$ of the first 100 LBO eigenvalues on two tetrahedral meshes, and then evaluate the difference $|\bm{\lambda}_{\mathcal{M}}-\bm{\lambda}_{\mathcal{N}}|$ Since the absolute difference between LBO eigenvalues is notoriously not informative, we compute both the relative difference and the offset difference proposed by Moschella et al. (2022). For each pair of tetrahedral meshes, we also extract the triangular meshes representing their boundary surface and compute the same quantities. The error distributions are shown at the bottom of Figure 2, where we can see that the spectral similarity between pairs of tetrahedral and surface meshes is comparable.
Volumetric Functional Maps.
Given a definition for the discrete LBO, the definition of the functional maps framework in tetrahedral meshes follows directly from its 2-dimensional counterpart (see Section 3). Nonetheless, in order to ensure a seamless integration of the existing approaches in the volumetric setting, we study how the usual properties of a functional map behave on tetrahedral meshes.
Given two meshes $\mathcal{M}$ and $\mathcal{N}$ and a functional map $\mathbf{C}:\mathcal{F}(\mathcal{M})\to\mathcal{F}(\mathcal{N})$, Lescoat et al. (2020) propose two metrics for evaluating the ideal qualities of $\mathbf{C}$: the orthogonality $\|\mathbf{C}\|_{O}$ and the Laplacian commutativity $\|\mathbf{C}\|_{L}$
(5) $\|\mathbf{C}\|_{O}=\frac{\|\mathbf{C}^{\top}\mathbf{C}-\mathbf{I}\|}{\|\mathbf{I}\|}\,,\qquad\|\mathbf{C}\|_{L}=\frac{\|\mathbf{C}\mathbf{\Lambda}_{\mathcal{M}}-\mathbf{\Lambda}_{\mathcal{N}}\mathbf{C}\|}{\|\mathbf{C}\|}\,,$
[…]
Figure 3. Distribution of the orthogonality (left) and Laplacian commutativity (right) metrics for pairs of volumetric shapes (blue) and surface meshes (orange).
Again, we compute the same quantities for tetrahedral and triangular meshes across the same 40 pairs of shapes, summarizing the results in Figure 3. As shown in the figure, the quality of volumetric functional maps is comparable to the surface cases.
4.2. Spectral Connectivity Transfer
The success of the functional maps framework is also due to its ability of smoothing out signals and correspondences. Indeed, the use of a truncated basis results in filtering out high-frequency details and producing a smoothly interpolated signal.
This smoothness property can be combined with the fact that the basis is usually much smaller than the number of vertices. As a result, the spectral representation of a signal can be approximated even if the signal is noisy or partially absent from some regions of the shape. This idea has been exploited by Melzi et al. (2019b) for efficiently approximating a functional map using only a reduced sampling of the original surface. In our setting, we take advantage of this property for transferring the connectivity between volumes.
Functional Maps for Connectivity Transfer.
Let $\mathcal{M}=(V_{\mathcal{M}},T_{\mathcal{M}})$ be a tetrahedral mesh and $\partial\mathcal{M}=(V_{\partial\mathcal{M}},F_{\partial\mathcal{M}})$ be its external surface. Let us consider another surface $\partial\mathcal{N}=(V_{\partial\mathcal{N}},F_{\partial\mathcal{N}})$ in known correspondence with $\partial\mathcal{M}$ through the (surface) mapping function $\pi:\partial\mathcal{M}\to\partial\mathcal{N}$. There are various options in the literature to obtain the tetrahedralization $\mathcal{N}$ of the interior of $\partial\mathcal{N}$ (Hang, 2015; Diazzi et al., 2023). However, by no mean the connectivity generated by these algorithms will match the connectivity of the mesh $\mathcal{M}$.
Given the surface $\partial\mathcal{N}$ and its induced tetrahedral mesh $\mathcal{N}$, we exploit the known correspondence $\pi:\partial\mathcal{M}\to\partial\mathcal{N}$ to induce a full correspondence $\pi^{\prime}:\mathcal{M}\to\mathcal{N}$. Let $\mathbf{\Phi}_{\mathcal{M}},\mathbf{\Phi}_{\mathcal{N}}$ be the LBO eigenbases of the volumes $\mathcal{M}$ and $\mathcal{N}$, respectively, and let us denote with $\mathbf{\Phi}_{\partial\mathcal{M}}=(\mathbf{\Phi}_{\mathcal{M}})_{\partial},\mathbf{\Phi}_{\partial\mathcal{N}}=(\mathbf{\Phi}_{\mathcal{N}})_{\partial}$ their restrictions to the surfaces $\partial\mathcal{M}$ and $\partial\mathcal{N}$, respectively. We stress that $\mathbf{\Phi}_{\partial\mathcal{M}}$ is not the LBO eigenbasis of $\partial\mathcal{M}$. We note that the functional map $\mathbf{C}$ induced by $\pi^{\prime}$ must be such that
(6) $\mathbf{\Phi}_{\mathcal{M}}\mathbf{C}=T_{\pi^{\prime}}(\mathbf{\Phi}_{\mathcal{N}})\,.$
We consider that the rows of $\mathbf{\Phi}_{\partial\mathcal{M}}$ (resp. $\mathbf{\Phi}_{\partial\mathcal{N}}$) are a subset of the rows of $\mathbf{\Phi}_{\mathcal{M}}$ (resp. $\mathbf{\Phi}_{\mathcal{N}})$. Recalling that $\pi$ is the restriction of $\pi^{\prime}$, we get
(7) $\mathbf{\Phi}_{\partial\mathcal{M}}\mathbf{C}=(\mathbf{\Phi}_{\mathcal{M}})_{\partial}\mathbf{C}=(\mathbf{\Phi}_{\mathcal{M}}\mathbf{C})_{\partial}=(T_{\pi^{\prime}}(\mathbf{\Phi}_{\mathcal{N}}))_{\partial}=T_{\pi}(\mathbf{\Phi}_{\partial\mathcal{N}})\,.$
The size of the basis (i.e., the number of columns of $\mathbf{\Phi}_{\partial\mathcal{M}}$) is generally smaller than the number of vertices (i.e., the number of rows), therefore $\mathbf{\Phi}_{\partial\mathcal{M}}$ admits left inverse and the volumetric functional map can be approximated via the known surface correspondence as $\mathbf{C}\approx\mathbf{\Phi}_{\partial\mathcal{M}}^{\dagger}T_{\pi}(\mathbf{\Phi}_{\partial\mathcal{N}})$.
The functional map $\mathbf{C}$ can then be used with the full spectrum to transfer the coordinates of $\mathcal{N}$ onto $\mathcal{M}$. Namely
(8) $\begin{gathered}T_{\pi^{\prime}}(x_{\mathcal{N}})=\mathbf{\Phi}_{\mathcal{M}}\mathbf{C}\mathbf{\Phi}_{\mathcal{N}}^{\dagger}x_{\mathcal{N}}\,,\\ T_{\pi^{\prime}}(y_{\mathcal{N}})=\mathbf{\Phi}_{\mathcal{M}}\mathbf{C}\mathbf{\Phi}_{\mathcal{N}}^{\dagger}y_{\mathcal{N}}\,,\\ T_{\pi^{\prime}}(z_{\mathcal{N}})=\mathbf{\Phi}_{\mathcal{M}}\mathbf{C}\mathbf{\Phi}_{\mathcal{N}}^{\dagger}z_{\mathcal{N}}\,.\end{gathered}$
By using the values $T_{\pi^{\prime}}(x_{\mathcal{N}}),T_{\pi^{\prime}}(y_{\mathcal{N}}),T_{\pi^{\prime}}(z_{\mathcal{N}})$ as coordinates for the vertices of $\mathcal{M}$, we get a connectivity transfer from $\mathcal{M}$ onto $\mathcal{N}$.
Spectral Representation of Volumes.
Intuitively, a watertight surface already encodes information about the enclosed volume. Since the LBO eigenbasis provides smoothness guarantees in signal reconstruction, it is a good candidate for extrapolating the interior coordinates of a volume from its surface.
Knowing the ground truth correspondence $\pi:\partial\mathcal{M}\to\partial\mathcal{N}$, we can transfer the surface coordinates $x_{\partial\mathcal{N}},y_{\partial\mathcal{N}},z_{\partial\mathcal{N}}$ of $\partial\mathcal{N}$ onto $\partial\mathcal{M}$ as $T_{\pi}(x_{\partial\mathcal{N}}),T_{\pi}(y_{\partial\mathcal{N}}),T_{\pi}(z_{\partial\mathcal{N}})$. We then consider the LBO eigenbasis $\mathbf{\Phi}_{\mathcal{M}}$ and its restriction $\mathbf{\Phi}_{\partial\mathcal{M}}$ to the surface. Following a reasoning analogous to the discussion in the previous paragraph, we have that
(9) $\begin{gathered}(T_{\pi^{\prime}}(x_{\mathcal{N}}))_{\partial}\approx(\mathbf{\Phi}_{\mathcal{M}}\mathbf{\Phi}_{\mathcal{M}}^{\dagger}T_{\pi^{\prime}}(x_{\mathcal{N}}))_{\partial}=\mathbf{\Phi}_{\partial\mathcal{M}}\mathbf{\Phi}_{\mathcal{M}}^{\dagger}T_{\pi^{\prime}}(x_{\mathcal{N}})\,,\\ T_{\pi}(x_{\partial\mathcal{N}})\approx\mathbf{\Phi}_{\partial\mathcal{M}}\mathbf{\Phi}_{\partial\mathcal{M}}^{\dagger}T_{\pi}(x_{\partial\mathcal{N}})\,.\end{gathered}$
Since $(T_{\pi^{\prime}}(x_{\mathcal{N}}))_{\partial}=T_{\pi}(x_{\partial\mathcal{N}})$, then it must also be $\mathbf{\Phi}_{\mathcal{M}}^{\dagger}T_{\pi^{\prime}}(x_{\mathcal{N}})\approx\mathbf{\Phi}_{\partial\mathcal{M}}^{\dagger}T_{\pi}(x_{\partial\mathcal{N}})$. Finally, the coordinates on the interior of $\mathcal{N}$ can be extrapolated by reconstructing with the full basis:
(10) $\begin{gathered}T_{\pi^{\prime}}(x_{\mathcal{N}})=\mathbf{\Phi}_{\mathcal{M}}\mathbf{\Phi}_{\partial\mathcal{M}}^{\dagger}T_{\pi}(x_{\partial\mathcal{N}})\,,\\ T_{\pi^{\prime}}(y_{\mathcal{N}})=\mathbf{\Phi}_{\mathcal{M}}\mathbf{\Phi}_{\partial\mathcal{M}}^{\dagger}T_{\pi}(y_{\partial\mathcal{N}})\,,\\ T_{\pi^{\prime}}(z_{\mathcal{N}})=\mathbf{\Phi}_{\mathcal{M}}\mathbf{\Phi}_{\partial\mathcal{M}}^{\dagger}T_{\pi}(z_{\partial\mathcal{N}})\,.\end{gathered}$
Concrete examples of using these two approaches for reliable connectivity transfer are depicted in Figure 8. For this specific test, we extended the LBO eigenbasis with the Coordinates Manifold Harmonics (CMH) basis (Marin et al., 2019), which also includes three additional functions encoding the extrinsic per vertex $xyz$ coordinates and is known to be better suited to reconstruct extrinsic information. Additional experiments on the whole dataset, considering both the LBO and CMH eigenbasis are reported in Section 5.2.
4.3. Volume-Aware Surface Correspondence
Besides purely volumetric applications, we observe that a good functional map among volumes also induces a good functional map among their external surfaces. Intuitively, a volumetric shape is a richer structure than a surface, yielding more information both locally and globally. Therefore, we guess that functional maps among tetrahedral meshes are more accurate and informative than maps among triangular meshes.
As mentioned in Section 4.2, the functional map that relates the LBO eigenbases of two volumetric shapes also relates their restriction to the surfaces. As a consequence, the volumetric map can be used to infer a correspondence between surfaces.
Given two surface meshes $\partial\mathcal{M},\partial\mathcal{N}$, we compute the tetrahedral meshes $\mathcal{M},\mathcal{N}$ representing their interior volume. In general, the tetrahedralization of a surface does not always produce a volume with the same surface connectivity as the original mesh, but if the two shapes can be closely aligned in 3D the scalar functions can be faithfully transferred using nearest surface point projections with barycentric interpolation (Maggioli et al., 2025).
Once we have the tetrahedral meshes $\mathcal{M},\mathcal{N}$, we can compute the functional map $\mathbf{C}$ relating their eigenbases via $\mathbf{\Phi}_{\mathcal{M}}\mathbf{C}=T_{\pi}(\mathbf{\Phi}_{\mathcal{N}})$, and we can use that same functional map for defining the surface relation $(\mathbf{\Phi}_{\mathcal{M}})_{\partial}\mathbf{C}=(T_{\pi}(\mathbf{\Phi}_{\mathcal{N}}))_{\partial}$. Therefore, the surface correspondence can be extracted via the nearest neighbor search $\mathrm{NN}(\mathbf{\Phi}_{\partial\mathcal{M}}\mathbf{C},\mathbf{\Phi}_{\partial\mathcal{N}})$.
5. Results and Applications
We implemented our software pipeline in Matlab, starting from the reference implementations for the functional map estimation with descriptors preservation (Nogneng and Ovsjanikov, 2017) and ZoomOut (Melzi et al., 2019b), extending these tools to enable volume spectral processing, as detailed in Section 4 and in the remainder of this section.
5.1. Volumetric Functional Maps
Figure 4. Left: average geodesic error curves on the pairs from Su et al. (Su et al., 2019) resulting from the application of FMaps (Nogneng and Ovsjanikov, 2017), ZoomOut (Melzi et al., 2019b), and Orthoprods (Maggioli et al., 2021). Right: The execution time for the three methods plotted against the number of vertices in the source mesh.
Figure 5. Coordinate transfer between two volumetric shapes using the Orthoprods basis. The normalized coordinates are treated as RGB channels for visualization (first column). For further validation, the transferred coordinates are also used for generating an error sensitive procedural texture, which is visualized on different slices of the volume (second to fourth columns).
Figure 6. Number of flipped tetrahedra obtained by transferring the mesh connectivity considering a growing number of basis functions. From left to right: coordinate transfer using the LBO basis; coordinate transfer using the CMH basis; coordinate extrapolation using the LBO basis; coordinate extrapolation using the CMH basis. The smaller colored lines show the trend for each pair of shapes in our reference dataset (Su et al., 2019). The bold black line is the average trend across the entire dataset.
To validate our prototype and assess its ability to devise a high-quality volume map, we compare the correspondences we compute against the ground truth. Unlike the surface setting, where in recent years there has been a constantly growing amount of public data useful for validation, almost no volumetric datasets are available for our specific task. We consider a set of 40 pairs of tetrahedral meshes with identical connectivity, which thus define an injective piece-wise linear map as described in the beginning of Section 2.1. This dataset, that was originally proposed by (Su et al., 2019) and then used to validate (Du et al., 2020), covers diverse classes of objects, such as: animals, humanoids, hands, and a variety of man-made objects like glasses, lamps, laptops and scissors. For each pair of objects we compute a volumetric functional map and extract a point-to-point correspondence as described in Section 4.1. Then, for each vertex $v$ we measure the geodesic error, which is the geodesic distance $d_{\mathcal{N}}(\pi_{\mathrm{gt}}(v),\pi(v))$ between its mapped point $\pi(v)$ and its ground truth counterpart $\pi_{\mathrm{gt}}(v)$. In Figure 4 we show the cumulative curves of the geodesic error obtained with different functional map estimation algorithms, as well as their runtime in relation to the vertex count. We also provide the average geodesic error (AGE) in the legend of the figure. For the comparison, we use the well-established algorithm based on descriptors preservation (FMap) (Nogneng and Ovsjanikov, 2017) and ZoomOut (Melzi et al., 2019b), using both the standard and the fast implementation. For the fast implementation, we use the surface points as samples to validate the claim in Section 4.2 that the surface encodes information about the enclosed volume. We also test our framework replacing the LBO eigenbasis with the orthogonalized eigenproducts basis (Orthoprods) proposed by Maggioli et al. (2021), but instead of using the approach described by the authors for the functional map estimation, we plug the basis into the ZoomOut pipeline. As expected, the extended Orthoprods basis has a higher descriptive power, but this comes with an additional computational cost for aligning the larger bases. In contrast, ZoomOut achieves the best trade-off between accuracy and runtime. Overall, our results demonstrate that the volumetric functional map framework produces high quality maps, matching the performance obtained in the surface setting. This can be further appreciated in the signal transfer example from Figure 5, where we transfer the coordinates function between two meshes with the Orthoprods bases and use the resulting values to compute an error sensitive procedural texture (Maggioli et al., 2022).
Method 5% eigs 10% eigs 15% eigs 20% eigs Time (s)
Transfer (LBO) 4.63% 1.50% 1.18% 0.93% 1127
Transfer (CMH) 4.75% 0.84% 0.45% 0.34% 1127
Extrapolation (LBO) 4.39% 1.45% 1.15% 0.92% 554
Extrapolation (CMH) 4.90% 0.99% 0.56% 0.42% 554
Table 1. Average percentage of flips for the methods from Figure 6 at varying of the basis size. The last column also report the average time (in seconds) required for the computation of the largest basis considered.
5.2. Connectivity Transfer
Representing volume maps using two tetrahedral meshes that share the same connectivity but have different vertex coordinates is a common choice for many practical applications involving digital fabrication (Liu et al., 2024; Zhang et al., 2022; Etienne et al., 2019), medicine (Abulnaga et al., 2021) and hexahedral meshing (Pietroni et al., 2022; Brückler and Campen, 2023). Our framework can be used to transfer the connectivity of a given tetrahedral mesh onto a target domain, either by using a volumetric functional map or by extrapolating the connectivity from its outer surface, as detailed in Section 4.2.
Figure 7. Left: one of the 1405 failure cases of the untangling algorithm proposed by Garanzha et al. (2021), which emerged from the challenging large scale validation described in (Nigolian et al., 2023, §6.3). Right: computing an initial connectivity transfer with our tool to warm start the same algorithm allows to retain a fully bijective map, thus improving the overall robustness.
Figure 8. From left to right: a tetrahedral mesh; the external surface of the target domain; the ground truth connectivity transfer, obtained with the volumetric ARAP implementation in Cinolib (Livesu, 2019); the connectivity transfer obtained by extrapolating the interior vertices from the surface coordinates; the connectivity transfer obtained by transferring the coordinates with the functional map. Each tetrahedron $t$ is colored according to the geometric distortion of the map, measured as $|1-\det(J_{t})|$, where $J_{t}$ is the Jacobian of the affine transformation mapping $t$ from the source to the target shape. Both the surface extrapolation and volumetric transfer approaches yield maps that are visually identical with the ground truth. All maps shown in this figure – both the ground truth and the ones that we computed – are fully injective (i.e., $\forall t,\det(J_{t}$)¿0).
If the whole spectrum is considered, the LBO basis spans the entire functional space of the mesh and the functional map yields one-to-one correspondences (Rodola et al., 2017). Computing the full spectrum is computationally expensive, hence only a fraction of the eigenfunctions are typically computed. We test our prototypes on the dataset of 40 meshes released by Su  et al. (2019) and empirically verify that the amount of inverted elements introduced in the piece-wise map gracefully decays to zero for growing numbers of eigenfunctions (Figure 6). On average, already using $5\%$ of the spectrum the number of flipped tetrahedra is below $5\%$. It goes below $1\%$ if $20\%$ of the spectrum is considered (Table 1). We consider both the LBO spectrum and its extension equipped with three additional functions that encode the extrinsic embedding (CMH). As shown in (Marin et al., 2019; Melzi et al., 2020) CMH is better suited to transfer discrete signals such as the surface mesh connectivity. This advantage also manifests on volumes, as reflected in the second and fourth columns in Figure 6 and in the second and fourth lines of Table 1. For efficiency, the average running times for computing the larger bases are reported in the last column of the same table. As can be noticed the extrapolation algorithm is roughly 2×2$\times$2 × faster than volumetric transfer, because in this case only the volumetric spectrum of a single mesh must be computed.
Our results on connectivity transfer suggest that, even when the computational resources are limited, our tool could be used to reliably warm start mesh untangling methods such as (Garanzha et al., 2021; Du et al., 2020), which will then be required to sanitize the map by removing only a fraction of flipped tetrahedra, improving the overall robustness (Figure 7). Interestingly, if the mesh resolution is sufficiently low and the computation of a higher percentage of the spectrum is computationally feasible, we can directly devise a fully injective piecewise linear map, as shown in Figure 8, where $25\%$ of the spectrum is considered.
Figure 9. A segmentation of a template brain transferred to two other brains using a correspondence computed with our volumetric functional maps framework. The shapes and the segmentation are from MedShapeNet (Li et al., 2024).
5.3. Segmentation Transfer
The functional maps framework proved to be particularly successful in the area of statistical shape analysis (Stegmann and Gomez, 2002). In particular, the study of medical data primarily benefits from accurate solutions for analyzing large collections of shapes (Heimann and Meinzer, 2009), and the literature proves that the functional maps framework can be successfully applied for medical applications (Melzi et al., 2016; Magnet et al., 2023). A standard approach for analyzing collections of shapes is to determine or build a template surface which acts as a median shape for comparing the entire dataset (Huang et al., 2019). However, existing approach are limited to the analysis of surfaces, while often medical data are better represented as 3D volumes, such in the case of MRI scans (Rinck, 2019).
In this setting, our approach allows to extend existing methodologies to deal with volumetric representations. To prove this, we select two brain shapes from the MedShapeNet dataset (Li et al., 2024) and a population-averaged template (Evans et al., 1993). Using TetGen (Hang, 2015), we compute the tetrahedral meshes representing the volumes of the brains, and segment the template using a reference segmentation (Fischl et al., 2002). We then compute a volumetric functional map from the template to each shape, extract the point-to-point correspondence and transfer the segmentation. The results are reported in Figure 9, where we show the segmentation on both the surface and on a slice of the volume. We can see that, despite the noisy segmentation in the central section of both the source and the target domains, our tool can faithfully transfer the segmentation preserving its overall structure.
5.4. Shape Matching
Besides the ability to transfer volumetric signals, we also study to what extent incorporating volumetric information onto the shape matching problem improves accuracy over existing surface-based functional maps pipelines.
Data.
We consider both the outer surfaces of our reference ground truth volumetric dataset (which we will refer to as VOL) (Su et al., 2019) and well established datasets for shape matching, such as SHREC’19 (Melzi et al., 2019a), SHREC’20 (Dyke et al., 2020), and TOPKIDS (Lähner et al., 2016), which we all tetrahedralize using fTetWild (Hu et al., 2020). Overall, these datasets stress shape matching algorithms in different ways, exhibiting geometric and topological noise (SHREC’19), strongly non-isometric matches (SHREC’20) and topological changes (SHREC’20 and TOPKIDS). For SHREC’20, we discard the two partial meshes shown in Figure 10 because the inside/outside system of fTetWild was not able to correctly reproduce a volumetric mesh that allows a meaningful comparison with surface methods. For the same reason, we are not able to consider alternative datasets such as SMAL (Zuffi et al., 2017), which contains numerous self-intersections that yield topological changes during tetrahedralization. Attempts to sanitize these models with various combinations of geometry processing tools (Portaneri et al., 2022; Hang, 2015; Livesu, 2019; Diazzi and Attene, 2021; Diazzi et al., 2023) ended with no luck. Indeed, our difficulties in creating a dataset that allows to meaningfully compare surface-based and volume-based shape matching methods raises an open research challenge that will hopefully be addressed by the community in the near future (Section 6).
Figure 10. The two partial surfaces from the SHREC’20 dataset that we discarded from our experiments because once turned into tetrahedral meshes they do not correctly convey the intended shape, introducing an unwanted bias in the comparative analysis with surface-based methods.
Figure 11. Cumulative plot of the geodesic error obtained with the standard functional maps framework (blue) on surfaces and our volumetric functional maps approach (orange) on four different datasets. The results are averaged across each dataset.
Dataset Basis AGE Surf. AGE Vol. Succ. Rate Slowdown
[…]
Table 2. The specific functional maps implementation used for each dataset. We also report the geodesic error averaged across the entire dataset (AGE) of both the surface and volumetric approach, as well as the percentage of single pairs where our approach obtains a better AGE (Succ. Rate) and the average slowdown factor induced by using our method. As can be noticed, computational overhead heavily correlates with mesh growth, which we measure as the average ratio between the number of vertices in the volume and the surface meshes and we report between parentheses.
Comparative Analysis.
For validating the approach discussed in Section 4.3, we compare it with the standard functional maps framework for surfaces on each dataset presented in the previous paragraph. In order to prove that our volumetric framework can be seamlessly integrated with existing extensions of functional maps, we also test it with different sets of bases, such as the Coordinates Manifold Harmonics (CMH) (Marin et al., 2019) and the orthogonalized eigenproducts (Orthoprods) (Maggioli et al., 2021). In all cases, we align the bases using the method based on descriptors preservation (Nogneng and Ovsjanikov, 2017) and ZoomOut (Melzi et al., 2019b). For a fair comparison, we use the same basis for both the surface and volumetric approach. Details can be found in Table 2
Figure 12. The tool we use to tetrahedralize surface matching datasets (Hu et al., 2020) is robust, but occasionally introduces topological noise, closing unwanted handles that alter the mapping task (closeup). The LBO eigenbasis is not robust against such defects. Conversely, the CMH basis is robuster and greatly improves the mapping quality. Both the camel and the cow are from the SHREC’20 dataset.
The table also reports the dataset averaged geodesic error (AGE) for both the frameworks and the percetange of single pairs for which our approach obtains a better AGE (Success Rate), as well as the slowdown factor induced by the introduction of the volume and the ratio between the vertex count of the tetrahedral and the surface meshes (between parentheses). More detailed results about the comparison can be found in Figure 11, where we show the cumulative geodesic error curves averaged across each dataset and their corresponding area under the curve (AUC). Interestingly, our approach is consistently much more accurate than its surface counterpart, suggesting that incorporating volumetric information into the matching process is always a rewarding choice when a volumetric discretization is available. The increase in accuracy comes at the cost of some additional computational effort, primarily due to the introduction of the internal vertices. By inspecting the last column of Table 2, we can see that there is a close relationship between the slowdown and the number of vertices introduced, suggesting that the volumetric pipeline is not significantly slower per se, but suffers from the overhead induced by the additional vertices needed to discretize the volume.
SHREC’20 is the only dataset where the surface-only method occasionally obtains better results (in the 7.27% of the cases). Based on our understanding, these sporadic cases in which surface-based methods surpass our volume pipeline depend on the basis alignment algorithm we use (Nogneng and Ovsjanikov, 2017), which is known to not be robust under strong non-isometric deformations and topological changes, which occasionally occurred when tetrahedralizing our surface-based reference datasets. This behavior can be partially compensated by using more informative bases such as CMH, which provide a better tool for representing extrinsic information. This can be better appreciated in the example from Figure 12, where we show a surface coordinate transfer in a pair from the SHREC’20 dataset with our volumetric approach using the standard LBO eigenbasis and the CMH basis. However, integrating more robust basis alignment methods is certainly worth of future investigation.
6. Conclusions
We have presented the first volumetric extension of the functional maps pipeline, showcasing practical applications that build on top of this connection to transfer both discrete and continuous signals for solid texturing, connectivity transfer and segmentation transfer for medical applications, as well as to improve the accuracy of shape matching tasks on surfaces. We believe that our practical contributions just scratch the surface of this problem and foresee interesting avenues for future research.
Limitations.
Our results on connectivity transfer show great potential but are far from being perfect. Indeed, achieving a fully injective map is technically possible, but considering a large enough portion of the spectrum may easily become computationally infeasible for applications. Devising smarter techniques to retain more precision for fixed spectrum size is an interesting direction worth investigating.
The higher precision in shape matching tasks comes at the cost of an additional computational effort, because the most expensive steps in the pipeline (i.e., the eigendecomposition and the alignment of the bases) suffer from the increased vertex count of the volume mesh. This limitation, which is intrinsic of our approach, could be mitigated by integrating scalable algorithms for computing and aligning the bases specifically designed for volume meshes.
Future works.
First and foremost, a theoretical study of what properties are invariant under a volume functional map must be carried out. Despite our empirical studies show that volumetric maps are more accurate than their surface counterparts, theoretically justifying this behavior would provide useful insights for the development of new solutions.
Secondly, the success of the functional maps framework was also related to its high flexibility and easy extensibility. A variety of solutions have been devised to extend the LBO eigenbasis, refine the map, and integrate the framework in machine learning pipelines. In this work, we just scratched the surface of this line of research, showing that some of the existing extensions can be seamlessly integrated into the volumetric framework. However, other works would require deeper analyses to generalize to tetrahedral meshes.
Finally, we emphasize the importance of releasing new volumetric datasets to support ongoing research on volume mapping in general, and on spectral volume methods in particular. As discussed in Section 5.4 surface-based datasets that are well established in the functional map community hardly extend to volumes because of the difficulty to produce tetrahedral meshes of acceptable quality, due to the presence of both topological and geometric noise, (e.g., large missing parts, overlaps, multiple connected components, non-manifold configurations and self-intersections). While robust volume meshing methods such as (Hu et al., 2020; Diazzi and Attene, 2021) can handle intersections with ease, determining semantically plausible separations between the inside and outside portions of the volume in the presence of large gaps remains an open meshing challenge (Figure 10). Recently, a large scale dataset to support ongoing research on piece-wise volume mappings to convex or star-shaped domains has been released (Cherchi and Livesu, 2023). Having similar contributions to support spectral methods would greatly help practitioners and ultimately foster more research in this emerging field.